# URLOST: Unsupervised Representation Learning without Stationarity or Topology

### v1

@@ 3.2 V1 neural response to natural image stimulus
After accessing URLOST’s performance on synthetic biological vision data, we take a step further to challenge its generalizability with high-dimensional natural datasets. The first task is decoding neural response recording in the primary visual area (V1) of mice.
V1 neural response dataset. The dataset, published by [44], contains responses from over 10,000 V1 neurons captured via two-photon calcium imaging. These neurons responded to 2,800 unique images from ImageNet [12], with each image presented twice to assess the consistency of the neural response. In the decoding task, a prediction is considered accurate if the neural response to a given stimulus in the first presentation closely matches the response to the same stimulus in the second presentation within the representation space. This task presents greater challenges than the synthetic biological vision described in the prior section. For one, the data comes from real-world neural recordings rather than a curated dataset like CIFAR-10. For another, the geometric structure of the V1 area is substantially more intricate than that of the retina. To date, no precise mathematical model of the V1 neural response has been well established. The inherent topology and stationarity of the data still remain difficult to grasp [43; 42]. Nevertheless, evidence of Retinotopy  [18; 19] and findings from prior research [41; 7; 58] suggest that the neuron population code in V1 are tiling a low dimensional manifold. This insight led us to treat the population neuron response as high-dimensional data and explore whether URLOST can effectively learn its representation.
Experiments. Following the approach in [44] we apply standardization and normalization to the neural firing rate. The processed signals are high-dimensional vectors, and they can be directly used for the decoding task, which serves as the “raw” signal baseline in Table 2. For representation learning methods, URLOST is evaluated along with MAE and $\beta$-VAE [26]. Note that the baseline methods need to handle high-dimensional vector data without stationarity or topology, so SimCLR is no longer applicable. We use $\beta$-VAE instead. We first train the neural network with an unsupervised learning task, then use the latent state of the network as the representation for the neural responses in the decoding task. The results are presented in the table 2. Our method surpasses the original neuron response and other methods, achieving the best performance.
@@ 3.3 Gene expression data
Gene expression dataset. The dataset comes from The Cancer Genome Atlas (TCGA) [62; 66], which is a project that catalogs the genetic mutations responsible for cancer using genome sequencing and bioinformatics. The project molecularly characterized over 20,000 primary cancers and matched normal samples spanning 33 cancer types. We focus on the pan-cancer classification task: diagnose and classify the type of cancer for a given patient based on his gene expression profile. The TCGA project collects the data of 11,000 patients and uses Micro-RNA (miRNA) as their gene expression profiles. Like the V1 response, no explicit topology and stationarity are known and each data point is a high-dimensional vector. Specifically, 1773 miRNA identifiers are used so that each data point is a 1773-dimensional vector. Types of cancer that each patient is diagnosed with serve as the classification labels.
Experiments. Similar to Section 3.2, URLOST is compared with the original signals, MAE, and $\beta$-VAE, which is the state-of-the-art unsupervised learning method on TCGA cancer classification [71; 72]. We also randomly partition the dataset do five-fold cross-validation and report the average performance in Table 2. Again, our method learns meaningful representation from the original signal. The learned representation benefited the classification task and achieved the best performance, demonstrating URLOST’s ability to learn meaningful representation of data from diverse domains.
@@ 4 Ablation study
(a) Replacing the non-shared projections of the self-organizing layer with the shared projection layer entails a significant drop in performance.
@@
Foveated SC 82.7 %
CIFAR-10 DSC 85.4 %
(b) “SC” denotes spectral clustering with uniform density clustering and “DSC” denotes density adjusted spectral clustering. For Foveated CIFAR-10 using density adjusted spectral clustering to create clusters will make the model perform better than using standard spectral clustering with uniform density.
4.1 Self-organizing Layer vs Shared Projection Layer
$z_{0}=[Ex^{(1)},\cdots Ex^{(M)}]$ (3)
which is further processed by a neural network. The sequential inputs can be a list of language tokens [14; 50], pixel values [5], image patches [15], or overlapped image patches [6; 24; 70]. $E$ can be considered as a projection layer that is shared among all elements in the input sequence. The self-organizing layer $g(\cdot,w^{(i)})$ introduced in Section 2.3 can be considered as a non-shared projection layer. We conducted an ablation study comparing the two designs to demonstrate the effectiveness of the self-organizing layers both quantitatively and qualitatively. To facilitate the ablation, we further synthesized another dataset.
Locally-permuted CIFAR-10. To directly evaluate the performance of the non-shared projection approach, we designed an experiment involving intentionally misaligned clusters. In this experiment, we divide each image into patches and locally permute all the patches. The $i$-th image patch is denoted by $x^{(i)}$, and its permuted version, permutated by the permutation matrix $E^{(i)}$, is expressed as $E^{(i)}x^{(i)}$. We refer to this manipulated dataset as the Locally-Permuted CIFAR-10. Our hypothesis posits that models using shared projections, as defined in Equation 3, will struggle to adapt to random permutations, whereas self-organizing layers equipped with non-shared projections can autonomously adapt to each patch’s permutation, resulting in robust performance. This hypothesis is evaluated quantitatively and through the visualization of learned weights $w^{(i)}$.
Quantitative results. Table 3 confirms our hypothesis, demonstrating a significant performance decline in models employing shared projections when exposed to permuted data. In contrast, the non-shared projection model maintains stable performance.
Visual evidence. Using linear layers to parameterize the self-organizing layers, i.e. let $g(x,W^{(i)})=W^{(i)}x$, we expect that if the projection layer effectively aligns the input sequence, $E^{(i)T}W^{(i)}$ should exhibit visual similarities. That is, after applying the inverse permutation $E^{(i)T}$, the learned projection matrix $W^{(i)}$ at each location should appear consistent or similar. The proof of this statement is provided in Appendix A.4. The model trained on Locally-Permuted CIFAR10 provides visual evidence supporting this claim. In Figure 4, the weights show similar patterns after reversing the permutations.
4.2 Density adjusted clustering vs Uniform Density Clustering
As explained in Section 2.2, the shape and size of each cluster depend on how the density function is defined. Let $q(i)$ represent the eccentricity, the distance from $i$th kernel to the center of the sampling lattice, and let $n(i)=\sum_{j}A_{ji}$ where $A$ is the affinity matrix, then the density is defined as:
$p(i)=q(i)^{-\alpha}n(i)^{\beta}.$ (4)
By setting $\alpha$ and $\beta$ nonzero, the density function is eccentricity-dependent. Setting both $\alpha$ and $\beta$ to zero will make $n(i)$ constant which recovers the uniform density spectral clustering. We vary the parameters $\alpha$ and $\beta$ to generate different sets of clusters for the foveated CIFAR-10 dataset and run URLOST using each of these sets of clusters. Results in Table 4 validate that the model performs better with density adjusted clustering. The intuitive explanation is that by adjusting the values of $\alpha$ and $\beta$, we can make each cluster carry similar amounts of information (refer to Appendix A.3.). A balanced distribution of information across clusters enhances the model’s ability to learn meaningful representations. Without this balance, masking a low-information cluster makes the prediction task trivial, while masking a high-information cluster will make the prediction task too difficult. In either scenario, the model’s ability to learn effective representations is compromised.
@@ 5 Additional related works
Topology in biological visual signal. 2-D topology of natural images is strong prior that requires many bits to encode [13; 2]. Such 2-D topology is encoded in the natural image statistic [55; 28], which can be recovered [53]. Optic and neural circuits in the retina result in a more irregular 2-D topology than the natural image, which can still be simulated [52; 46; 47; 45; 61; 30]. This information is further processed by the primary visual cortex. Evidence of retinotopy suggests the low-dimensional geometry of visual input from retina is encoded by the neuron in primary visual cortex [40; 20; 27; 19; 65; 48]. These evidences suggest we can recover the topology using signal from retinal ganglion cell and V1 neurons.
@@
Evidence of self-organizing mechanism in the brain. In computational neuroscience, many works use the self-organizing maps (SOM) as a computational model for V1 functional organization: [16; 60; 1; 17; 39; 31]. In other words, this idea of self-organizing is likely a principle governing how the brain performs computations. Even though V1 functional organizations are present at birth, numerous studies also indicate that the brain’s self-organizing mechanisms continue after full development [22; 54; 29].
Learning with signal on non-euclidean geometry. In recent years, researchers from the machine learning community have made efforts to consider geometries and special structures beyond classic images, text, and feature vectors. [33] treats an image as a set of points but depends on the 2D coordinates. The geometric deep learning community tries to generalize convolution neural networks beyond the Euclidean domain [3; 37; 11; 21]. Recent research also explores adapting the Transformer to domains beyond Euclidean spaces [10; 9]. However, none of them has tried to tackle the issue when the data has no explicit topology or stationarity, which is the focus of URLOST.
@@
Self-supervised learning. Self-supervised learning (SSL) has made substantial progress in recent years. Different SSL method is designed for each modality, for example: predicting the masked/next token in NLP[14; 50; 4], solving pre-text tasks, predicting masked patches, or building contrastive image pairs in computer vision [36; 25; 68; 6; 24; 70]. These SSL methods have demonstrated descent scalability with a vast amount of unlabeled data and have shown their power by achieving performance on par with or even surpassing supervised methods. They have also exhibited huge potential in cross-modal learning, such as the CLIP by [51]. However, we argue that these SSL methods are all built upon specific modalities with explicit topology and stationarity which URLOST goes beyond.
6 Discussion
@@
The success of most current state-of-the-art self-supervised representation learning methods relies on the assumption that the data has known stationarity and domain topology, such as the grid-like RGB images and time sequences. However, biological vision systems have evolved to deal with signals with less regularity. In this work, we explore unsupervised representation learning under a more general assumption, where the stationarity and topology of the data are unknown to the machine learning model and its designers. We argue that this is a general and realistic assumption for high-dimensional data in modalities of natural science. We propose a novel unsupervised representation learning method that works under this assumption and demonstrates our method’s effectiveness and generality on a synthetic biological vision dataset and two datasets from natural science that have diverse modalities. We also perform a step-by-step ablation study to show the effectiveness of the novel components in our model.
During experiments, we found that density adjusted spectral clustering is crucial for the quality of representation learning. How to adjust the density and obtain a balanced clustering for any given data or even learning the clusters end-to-end with the representation via back-propagation is worth future investigation. Moreover, our current self-organizing layer is still simple though it shows effective performance. Extending it to a more sophisticated design and potentially incorporating it with various neural network architectures is also worth future exploration.
In summary, our method offers a handy and general unsupervised learning tool when dealing with high-dimensional data of arbitrary modality with unknown stationarity and topology, particularly common in the field of natural sciences, where many present strong unsupervised learning baselines cannot directly adapt. We hope it can provide inspiration for work in related fields.

### v2

@@ 3.2 V1 Neural Response to Natural Image Stimulus
After testing URLOST’s performance on synthetic biological vision data, we take a step further to challenge its generalizability with high-dimensional natural datasets. The first task is decoding neural response recording in the primary visual area (V1) of mice.
V1 neural response dataset. The dataset, published by [57], contains responses from over 10,000 V1 neurons captured via two-photon calcium imaging. These neurons responded to 2,800 unique images from ImageNet [19], with each image presented twice to assess the consistency of the neural response. In the decoding task, a prediction is considered accurate if the neural response to a given stimulus in the first presentation closely matches the response to the same stimulus in the second presentation within the representation space. This task presents greater challenges than the synthetic biological vision described in the prior section. For one, the data comes from real-world neural recordings rather than a curated dataset like CIFAR-10. For another, the geometric structure of the V1 area is substantially more intricate than that of the retina. To date, no precise mathematical model of the V1 neural response has been well established. The inherent topology and stationarity of the data still remain difficult to grasp [56; 55]. Nevertheless, evidence of retinotopy  [25; 26] and findings from prior research [54; 13; 74] suggest that the neuron population code in V1 are tiling a low dimensional manifold. This insight led us to treat the population neuron response as high-dimensional data and explore whether URLOST can effectively learn its representation.
Experiments. Following the approach in Pachitariu et al. [57] we apply standardization and normalization to the neural firing rate. The processed signals are high-dimensional vectors, and they can be directly used for the decoding task, which serves as the “raw” signal baseline in Table 2. For representation learning methods, URLOST is evaluated along with MAE and $\beta$-VAE [35]. For MAE baseline, we obtain patches by randomly selecting different dimensions from the signal. Note that the baseline methods need to handle high-dimensional vector data without stationarity or topology. Since SimCLR and other constrastive learning model leverage these two properties to make positive pair, they are no longer applicable. We use $\beta$-VAE as a baseline instead. We first train the neural network with an unsupervised learning task, then use the latent state of the network as the representation for the neural responses in the decoding task. The results are presented in table 2. Our method surpasses the original neuron response and other methods, achieving the best performance.
@@ 3.3 Gene Expression Data
Gene expression dataset. The dataset comes from The Cancer Genome Atlas (TCGA) [79; 84], which is a project that catalogs the genetic mutations responsible for cancer using genome sequencing and bioinformatics. The project molecularly characterized over 20,000 primary cancers and matched normal samples spanning 33 cancer types. We focus on the pan-cancer classification task: diagnose and classify the type of cancer for a given patient based on his gene expression profile. The TCGA project collects the data of 11,000 patients and uses Micro-RNA (miRNA) as their gene expression profiles. Like the V1 response, no explicit topology and stationarity are known and each data point is a high-dimensional vector. Specifically, 1773 miRNA identifiers are used so that each data point is a 1773-dimensional vector. Types of cancer that each patient is diagnosed with serve as the classification labels.
Experiments. Similar to Section 3.2, URLOST is compared with the original signals, MAE, and $\beta$-VAE, which is the state-of-the-art unsupervised learning method on TCGA cancer classification [91; 92]. We also randomly partition the dataset do five-fold cross-validation and report the average performance in Table 2. Again, our method learns meaningful representation from the original signal. The learned representation benefited the classification task and achieved the best performance, demonstrating URLOST’s ability to learn meaningful representation of data from diverse domains.
@@ 4 Ablation study
Masking Unit Permuted Cifar10 Foveated Cifar10 TCGA Gene V1 Response
Clusters (URLOST) 86.4% 85.4% 94.9% 78.8%
Random patch 55.7% 51.1% 91.7% 73.9%
Individual dimension 56.7% 48.5% 88.3% 64.8%
Table 3: Performance of different datasets using different masking units. “Clusters (URLOST)” refers to the clusters formed using pairwise mutual information and spectral clustering. “Random patch” refers to patches formed by aggregating random dimensions. “Individual dimension” refers to using individual dimensions as masking units without any aggregation. The model is trained for the mask prediction task, and the probing accuracy is reported in the table.
4.1 Aggregating similar dimensions for masking unit
Clustering is at the heart of the proposed method. Why should similar dimensions be aggregated to form a masking unit? Intuitively, we assume that similar dimensions are used to sample similar regions of the underlying signals. Instead of masking each dimension, making the entire cluster force the model to learn high-level structure from the signal, allowing the model to learn a rich representation. For example, for the CIFAR10 example, similar pixels tend to sample similar regions, which together form image patches. For V1 data, each neuron encodes a simple pattern like an oriented edge, but neurons together code patterns like shape and contours. We provide experiments for using different masking units, as shown in table 3, aggregating similar dimensions significantly improves the performance.
4.2 Self-organizing Layer vs Shared Projection Layer
$z_{0}=[Ex^{(1)},\cdots Ex^{(M)}]$ (3)
which is further processed by a neural network. The sequential inputs can be a list of language tokens [21; 63], pixel values [11], image patches [22], or overlapped image patches [12; 33; 90]. $E$ can be considered as a projection layer that is shared among all elements in the input sequence. The self-organizing layer $g(\cdot,w^{(i)})$ introduced in Section 2.3 can be considered as a non-shared projection layer. We conducted an ablation study comparing the two designs to demonstrate the effectiveness of the self-organizing layers both quantitatively and qualitatively. To facilitate the ablation, we further synthesized another dataset.
Locally-permuted CIFAR-10. To directly evaluate the performance of the non-shared projection approach, we designed an experiment involving intentionally misaligned clusters. In this experiment, we divide each image into patches and locally permute all the patches. The $i$-th image patch is denoted by $x^{(i)}$, and its permuted version, permutated by the permutation matrix $E^{(i)}$, is expressed as $E^{(i)}x^{(i)}$. We refer to this synthetic dataset as the Locally-Permuted CIFAR-10. Our hypothesis posits that models using shared projections, as defined in Equation 3, will struggle to adapt to random permutations, whereas self-organizing layers equipped with non-shared projections can autonomously adapt to each patch’s permutation, resulting in robust performance. This hypothesis is evaluated quantitatively and through the visualization of the learned weights $w^{(i)}$.
Quantitative results. Table 4 confirms our hypothesis, demonstrating a significant performance decline in models employing shared projections when exposed to permuted data. In contrast, the model with self-organizing layer maintains stable performance.
Visual evidence. Using linear layers to parameterize the self-organizing layers, i.e. let $g(x,W^{(i)})=W^{(i)}x$, we expect that if the projection layer effectively aligns the input sequence, $E^{(i)T}W^{(i)}$ should exhibit visual similarities. That is, after applying the inverse permutation $E^{(i)T}$, the learned projection matrix $W^{(i)}$ at each location should appear consistent or similar. The proof of this statement is provided in Appendix A.5. The model trained on Locally-Permuted CIFAR10 provides visual evidence supporting this claim. In Figure 4, the weights show similar patterns after reversing the permutations. These observations demonstrate that URLOST can also be used as an unsupervised learning method to recover topology and enforce stationary on the signal. Note that Figure 4 shows that the self-organizing layer learns to “undo” the permutation. Additionally, the self-organizing layer also does some extra regular transformation to each patch. It is likely that the self-organizing layer learns to encode position information as transformations of some simple group structures.
4.3 Density adjusted clustering vs Uniform Density Clustering
As explained in Section 2.2, the shape and size of each cluster depend on how the density function $p(i)$ is defined as:
$p(i)=q(i)^{\alpha}n(i)^{-\beta}$ (4)
where $n(i)=\sum_{j\in\text{Top}_{k}(A_{ji})}A_{ji}$, $A$ is the affinity matrix, $q(i)$ represent the eccentricity, the distance from $i$th kernel to the center of the sampling lattice. Setting $\alpha$ and $\beta$ nonzero, the density function is eccentricity-dependent. Setting both $\alpha$ and $\beta$ to zero will make $n(i)$ constant which recovers the uniform density spectral clustering. We vary the parameters $\alpha$ and $\beta$ to generate different sets of clusters for the foveated CIFAR-10 dataset and run URLOST using each of these sets of clusters. Results in Table 5 validate that the model performs better with density-adjusted clustering. The intuitive explanation is that by adjusting the values of $\alpha$ and $\beta$, we can make each cluster carry similar amounts of information (refer to Appendix A.4.). A balanced distribution of information across clusters enhances the model’s ability to learn meaningful representations. Without this balance, masking a low-information cluster makes the prediction task trivial, while masking a high-information cluster will make the prediction task too difficult. In either scenario, the model’s ability to learn effective representations is compromised.
@@ 4.3 Density adjusted clustering vs Uniform Density Clustering
(a) Replacing the self-organizing layer with the shared projection layer entails a significant drop in performance.
@@
(b) “SC” denotes spectral clustering with uniform density clustering and “DSC” denotes density adjusted spectral clustering. Using DSC to create clusters outperform model without DSC.
@@
Foveated CIFAR-10 SC 82.7 %
DSC 85.4 %
@@ 5 Additional related works
Topology in biological visual signal. 2-D topology of natural images is a strong prior that requires many bits to encode [20; 3]. Such 2-D topology is encoded in the natural image statistic [71; 37]. Optic and neural circuits in the retina result in a more irregular 2-D topology than the natural image, which can still be simulated [66; 59; 60; 58; 78; 41]. This information is further processed by the primary visual cortex. Evidence of retinotopy suggests the low-dimensional geometry of visual input from retina is encoded by the neuron in primary visual cortex [53; 27; 36; 26; 83; 61]. These study provide evidence that topology under retinal ganglion cell and V1 neurons can be recovered. The theory and computational model of how visual system code encodes such 2-D is well-studied in computational neuroscience. The self-organizing map (SOM) was proposed as the first computational model by Kohonen in 1982. The algorithm produces a low-dimensional representation of a higher-dimensional dataset while preserving the topological structure of the data [43]. SOM is also motivated to solve the “unscramble” pixels problem by “descrambling” by mapping pixels into a 2D index set by leveraging the mutual information between pixels at different locations. More detail is in Appendix A.12. [67] tackles the same problem with manifold learning.
@@
Evidence of self-organizing mechanism in the brain. In neuroscience, many works use the self-organizing maps (SOM) as a computational model for V1 functional organization [23; 76; 1; 24; 52; 43]. In other words, this idea of self-organizing is a principle governing how the brain performs computations. Even though V1 functional organizations are present at birth, numerous studies indicate that the brain’s self-organizing mechanisms continue after full development [30; 68; 40].
Learning with signal on non-euclidean geometry. In recent years, researchers from the machine learning community have made efforts to consider geometries and special structures beyond classic images, text, and feature vectors. This is the key motivation for geometric deep learning and graph neural networks (GNN). Many works generalizes common operator for processing Euclidean data like 2d convolution and attention [7; 49; 18; 28]. Due to the natural of graph neural network, they often only work on limited data regime and do not scale to large data [69; 10; 9]. However, recent development in [32] shows this direction is prominent. They successfully scales GNN to ImageNet. We compared their proposed neural network architecture with the ViT backbone used in URLOST. Recent research also explores adapting the Transformer to domains beyond Euclidean spaces such as [16; 15; 29]. Ma et al. [45] treats an image as a set of points but relies on 2D coordinates. URLOST employs a single mutual information graph to define the topology of the high-dimensional signal. Gao et al. [29], on the other hand, is designed to handle graph data, where each data point corresponds to a distinct graph. It segments a graph into “subgraphs,” processes them with a GNN, and then passes the output to a transformer. This approach is undeniably more flexible but requires all subgraphs to be globally aligned. Furthermore, the self-organizing layer in URLOST generalizes the “patch resizer” mechanism from FlexiViT used in Beyer et al. [4]. Finally, the lines of works follows the “perceiver” architecture is very related to URLOST [39; 38]. “Perceiver” also implicitly aggregates similar dimensions of high dimensional data and processes the similar dimension with an attention mechanism. [88] is a follow-up work that combines the Perceiver architecture and masked-prediction objective. However, the key difference is that the ”perceiver” computes similarity between dimensions on a single example signal. For image data, this is essentially clustering pixels based on the pixel intensity of color in a signal image. On the other hand, URLOST computes similarity over statistics of the distribution of the signal. Additionally, ”perceiver” implicitly aggregates similar dimensions, while URLOST explicitly aggregates similar dimensions and recovers the original topology of the signal.
@@
Self-supervised learning. Self-supervised learning (SSL) has made substantial progress in recent years. Different SSL method is designed for each modality, for example: predicting the masked/next token in NLP[21; 63; 8], solving pre-text tasks, predicting masked patches, or building contrastive image pairs in computer vision [48; 34; 87; 12; 33; 90]. These SSL methods have demonstrated descent scalability with a vast amount of unlabeled data and have shown their power by achieving performance on par with or even surpassing supervised methods. They have also exhibited huge potential in cross-modal learning, such as the CLIP by Radford et al. [64].
6 Discussion, Limitations, and future direction
@@
The success of most current state-of-the-art self-supervised representation learning methods relies on the assumption that the data has known stationarity and domain topology. In this work, we explore unsupervised representation learning under a more general assumption, where the stationarity and topology of the data are unknown to the machine learning model and its designers. We argue that this is a general and realistic assumption for high-dimensional data in modalities of natural science. We propose a novel unsupervised representation learning method that works under this assumption and demonstrates our method’s effectiveness and generality on a synthetic biological vision dataset and two datasets from natural science that have diverse modalities. We also perform a step-by-step ablation study to show the effectiveness of the novel components in our model.
Unlike the self-organizing layer, the procedure for aggregating similar dimensions together is separate from the training of MAE. Given the effectiveness of the self-organizing layer, this design could be sub-optimal and could be improved. Learning the clusters end-to-end with the representation via back-propagation is worth investigating in the future. Additionally, the computational costs of computing pairwise mutual information and clustering both scales quadratically with the number of dimensions of the signal. Although each procedure only needs to be performed once, the computation could become too slow and infeasible for extremely high-dimensional data ($d>10,000$) with our current implementation as shown in Appendix A.9. In the Appendix, we provide a benchmark of the runtime of our implementation. Nevertheless, we think this problem could be resolved by a GPU implementation for clustering and mutual information. Additionally, although natural science datasets such as TCGA and V1 calcium imaging are considered large datasets in their respective domains, they are still small compared to the datasets used in computer vision. We expect larger natural science datasets to emerge as measuring and imaging technology advances. Additionally, adapting URLOST to support contrastive learning objectives, such as SimCLR, presents another intriguing direction.
7 Acknowledgement
We would like to specially thank Bruno Olshausen for suggesting key prior work on computational models and biological evidence for recovering the 2D topology of natural images. We also appreciate the helpful suggestions from Surya Ganguli and Atsu Kotani.